# Vivid Emotions

### v3

@@ 4.1 Omni-modal Text-Centric Alignment
Table 2: Comparison on vision-language and speech benchmarks. 1) EMOVA outperforms GPT-4o/4V and Gemini Pro 1.5 on 11 of the 14 vision-language benchmarks, providing a powerful open-sourced alternative. 2) Meanwhile, our EMOVA achieves state-of-the-art performance on Librispeech, surpassing its speech and omni-modal counterparts significantly. : reported by [43].
@@ 5.2 Comparison to SOTA Models
As an omni-modal model, EMOVA obtains comparable performance with the state-of-the-art VLLMs on multiple vision-language benchmarks, while showing superior proficiency in solving math problems needing precise visual content interpretation. Our EMOVA-7B surpasses GPT-4V by +7.3 on MathVerse, and our EMOVA-72B exceeds GPT-4o by +6.1 on MathVista. On 11 out of the 14 benchmarks, our EMOVA outperforms both GPT-4o/4V and Gemini Pro 1.5, providing a powerful open-sourced alternative.
@@ Acknowledgments.
We gratefully acknowledge supports of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This research has been made possible by funding support from the Research Grants Council of Hong Kong through the Research Impact Fund project R6003-21.
@@ Appendix C More on Benchmark Evaluation
To thoroughly evaluate our model’s vision-language abilities, 14 benchmarks covering four different aspects of the real-life scenarios are utilized for a comprehensive assessment across multiple domains. Moreover, Automatic Speech Recognition (ASR) and Text-to-speech (TTS) are adopted to evaluate speech-language abilities.
@@
ScienceQA-Img [56] and AI2D [33] are utilized to assess the model’s ability to deal with scientific questions and diagrams, which requires reasoning across various subjects and interpreting structured visual content.

### v4

@@ 4.1 Omni-modal Text-Centric Alignment
Table 2: Comparison on vision-language and speech benchmarks. 1) EMOVA outperforms GPT-4o/4V and Gemini Pro 1.5 on 11 of the 15 vision-language benchmarks, providing a powerful open-sourced alternative. 2) Meanwhile, our EMOVA achieves state-of-the-art performance on Librispeech, surpassing its speech and omni-modal counterparts significantly. : reported by [43].
@@ 5.2 Comparison to SOTA Models
As an omni-modal model, EMOVA obtains comparable performance with the state-of-the-art VLLMs on multiple vision-language benchmarks, while showing superior proficiency in solving math problems needing precise visual content interpretation. Our EMOVA-7B surpasses GPT-4V by +7.3 on MathVerse, and our EMOVA-72B exceeds GPT-4o by +6.1 on MathVista. On 11 out of the 15 benchmarks, our EMOVA outperforms both GPT-4o/4V and Gemini Pro 1.5, providing a powerful open-sourced alternative.
@@ Acknowledgments.
We gratefully acknowledge supports of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This work has been made possible by a Research Impact Fund project (RIF R6003-21) and a General Research Fund project (GRF 16203224) funded by the Research Grants Council (RGC) of the Hong Kong Government. This work was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region (Grants C7004-22G-1 and 16202523) and the Joint Centre for Artificial Intelligence (Grant FB453). This work is supported by National Key Research and Development Program of China (2024YFE0203100) and National Natural Science Foundation of China (No. 62441615, 62201484 and 62136005).
@@ Appendix C More on Benchmark Evaluation
To thoroughly evaluate our model’s vision-language abilities, 15 benchmarks covering four different aspects of the real-life scenarios are utilized for a comprehensive assessment across multiple domains. Moreover, Automatic Speech Recognition (ASR) and Text-to-speech (TTS) are adopted to evaluate speech-language abilities.
@@
MMMU [91], ScienceQA-Img [56] and AI2D [33] are used to assess models’ ability to deal with scientific questions and diagrams, which requires reasoning across various subjects and interpreting structured visual content.